Calculate the number of atoms in 58.8 dm³ of xenon gas at stp



1
Expert's answer
2022-02-10T03:19:05-0500

1 mole of matter contains 6.022×1023 elementary entities of the substance.

n (Xe) = V / 22.4 = 58.8 / 22.4 = 2.625 mol

N (Xe) = 2.625 x 6.022×1023 = 15.8×1023